ROTATIONAL SPECTROSCOPY OF HD^{18}O

\underline{JOHN C. PEARSON}^{a}, SHANSHAN YU, HARSHAL GUPTA and BRIAN J. DROUIN, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109.

The recent detection of HD^{18}O in Orion KL^{b} by the Herschel Space Observatory provided a strong motivation to revisit the rotational spectrum in order to obtain more accurate calculations of transition frequencies. Rotational ransitions were recorded in the 300 − 2700 GHz frequency range. Analysis of the combined microwave and infrared data sets with an Euler series Hamiltonian^{c} has facilitated determination of a set of precise rotational constants to support precision velocity measurements. The new rotational data also provides a means of evaluating the performance of the MARVEL algorithm used in the recent review of all available HDO data^{d}.

aA part of this work was performed at the Jet Propulsion Laboratory, California Institute of Technology under contract with the National Aeronautics and Space Administration. Copyright 2010\begin{matrix}c \\ ◯ \end{matrix} California Institute of Technology. All rights reserved.

bE. A. Bergin et al., Astron. Astrophys. \underline{521}, (L20), 2010.

cH. M. Pickett, J. C. Pearson, and C. E. Miller J. Mol. Spectrosc. \underline{233} (174), 2005.

dJ. Tennyson et al. J. Quant. Spectrosc. Radiat. Transfer \underline{111}(2160), 2010.